# A Machine Learning study of the two-dimensional antiferromagnetic $q$-state Potts model on the square lattice

### v1

@@ I Introduction
Recently, the ideas of Machine Learning techniques have broadened their applications impressively. The usage of these advanced methods is found those days in basically every fields of science, from fundamental research to industry applications. One such scientific field is the physics. In particular, neural networks (NN) are employed to study critical phenomena of many physical models Tan16 ; Car16 ; Nie16 ; Kim18 ; Li18 ; Chn18 ; Car19 ; Meh19 ; Don19 ; Fuk21 ; Che23 ; Suk24 ; Che25 .
@@
Instead of using real spin configurations to train a NN, here we follow a similar idea of Ref. Tan20.1 to conduct the training procedure. Specifically, on a square area, two artificially made configurations for which the associated spins have their values in 1 and -1 alternatively in both the rows and columns are used as the building unit of training set. This method of constructing the training set is used in Ref. Li25 to study the critical behaviors of 2D antiferromagnetic Ising model with multi-interactions on the square lattice.
@@ IV Numerical Results
To examine the critical behaviors of the studied models using the NN approach, we have conducted large scale Monte Carlo calculations (MC) using the Swendsen-Wang-Kotecky algorithm Wan89 ; Wan90 for $q=2,3,4,5,$ and 6. For every considered $L$, the simulated spin configurations at many various temperatures $T$ are recorded. In the following subsections, the MLP outcomes will be presented.
@@ IV.2 The NN outcomes related to $Q=3,4,5,6$
Figure 8: $R$ as functions of $T$ and $L=64$ and $L_{1}=32$ for $q=3,4,5$ and 6 models.
@@ V Discussions and Conclusions
In this study, we explore the critical behaviors of 2D antiferromagnetic $q$-state Potts models for $q=2,3,4,5$ and 6 on the square lattice. It is well-known that the $q=3$ model is ordered only at zero temperature and $q=4,5,6$ models are always disordered at any temperature.
@@
The simple MLP with the training strategy considered here has been applied successfully to study the critical phenomena of 2D antiferromagnetic Ising model with nearest and next-to-nearest interactions. It is demonstrated convincingly that these investigated phase transitions are first order and the signals for the two-peak structure in the probability distributions of the relevant quantities are much stronger than that related to the Monte Carlo method.
@@
The curves of $R$ and STD of $R$ corresponding to the 8-state ferromagnetic Potts model are shown as the left and the right panel of fig. 10, respectively. Apparently, sudden jumps occur exactly at the theoretical $T_{c}$. In addition, the large gaps shown in both panels of the figure suggests convincingly that the phase transition for the 8-state ferromagnetic Potts model is first order. This agrees with the expectation. While further studies are needed, the outcomes demonstrated in figs. 9 and 10 are encouraged.

### v2

@@ I Introduction
Recently, the ideas of Machine Learning techniques have broadened their applications impressively. The usage of these advanced methods is found those days in basically every fields of science, from fundamental research to industry applications. One such scientific field is the physics. In particular, neural networks (NN) are employed to study critical phenomena of many physical models Tan16 ; Car16 ; Nie16 ; Kim18 ; Li18 ; Chn18 ; Car19 ; Meh19 ; Don19 ; Fuk21 ; Tir22 ; Che23 ; Suk24 ; Che25 .
@@
Instead of using real spin configurations to train a NN, here we follow a similar idea of Ref. Tan20.1 to conduct the training procedure. Specifically, on a square area, two artificially made configurations for which the associated spins have their values in 1 and -1 alternatively in both the rows and columns are used as the building unit of training set. This method of constructing the training set is used in Ref. Li25 to study the critical behaviors of 2D antiferromagnetic Ising model with multi-interactions on the triangular lattice.
@@ IV Numerical Results
To examine the critical behaviors of the studied models using the NN approach described here, we have conducted large scale Monte Carlo calculations (MC) using the Swendsen-Wang-Kotecky algorithm Wan89 ; Wan90 for $q=2,3,4,5,$ and 6. For every considered $L$, the simulated spin configurations at many various temperatures $T$ are recorded. In the following subsections, the MLP outcomes will be presented.
@@ IV.2 The NN outcomes related to $Q=3,4,5,6$
Figure 8: $R$ as functions of $T$ for $L=64$ and $L_{1}=32$ for $q=3,4,5$ and 6 models.
@@ V Discussions and Conclusions
In this study, we explore the critical behaviors of 2D antiferromagnetic $q$-state Potts models for $q=2,3,4,5$ and 6 on the square lattice. It is well-known that the $q=3$ model is ordered only at zero temperature and $q=4,5,6$ models are always disordered at any (finite) temperature.
@@
The simple MLP with the training strategy considered here has been applied successfully to study the critical phenomena of 2D antiferromagnetic Ising model with nearest and next-to-nearest interactions. It is demonstrated convincingly that these investigated phase transitions are first order and the associated signals are very persuasive.
@@
The curves of $R$ and STD of $R$ corresponding to the 8-state ferromagnetic Potts model are shown as the left and the right panel of fig. 10, respectively. Apparently, sudden jumps occur exactly at the theoretical $T_{c}$. In addition, the large gaps shown in both panels of the figure suggests convincingly that the phase transition for the 8-state ferromagnetic Potts model is first order. This agrees with the expectation. It should be pointed out that the outcomes given in figs. 9 and 10 are obtained with $L_{1}=32$. When $L_{1}=64$ is used, only the STD of $R$ reveals the right critical information for these two models. It is likely that STD of $R$ is a more appropriate quantity than $R$ when ferromagnetic systems are studied. In any case, while further studies are needed, the outcomes demonstrated in figs. 9 and 10 are encouraged.
@@
Finally, it should be pointed out that the criterions for determining whether a phase transition is first order for the cases of $L=L_{1}$ and $L>L_{1}$ found in (revised version of) Ref. Li25 are different. In particular, the NN parameter batch_size used for $L=L_{1}$ and $L>L_{1}$ are 80 and 40, respectively. Interestingly, if one considers batch_size =40=40 for the case of $L=L_{1}$, then the resulting criterion is the same as the one of $L>L_{1}$. Both criterions associated with $L=L_{1}$ identify the phase transitions studied in Ref. Li25 correctly (First-order phase transitions). It may be interesting to explore this further to see if there is any subtlety when the configurations of fig. 2 are used as the training set(s).
@@ Acknowledgment
Conflict of Interest
The authors declare no conflict of interest.
Data Availability Statement
Data are available from the corresponding author on reasonable request.